OpenStudy (anonymous):

Which of the following equations is the equation 2 - 3y + 6x = 0 written in slope-intercept form? can anyone help and explain your steps? @iGreen @Kemist

OpenStudy (anonymous):

y = -2x - 2/3 y = 1/2x + 2 y = 2x + 2/3

OpenStudy (anonymous):

its one of those but idk how to do this one

OpenStudy (igreen):

Answer is the last one. y = 2x + 2/3

OpenStudy (anonymous):

thanks but can you explain how?

OpenStudy (igreen):

I just graphed it.. https://www.desmos.com/calculator Type each equation in a separate box to the left. The one that matches up with the one it tells you in the question is your answer.